# GraphLand: Evaluating Graph Machine Learning Models on Diverse Industrial Data

### v3

@@ B.1 Models
The models discussed above are graph-agnostic. To see if they can be improved by being granted access to some of the graph information, we augment them with neighborhood feature aggregation (NFA) — a simple feature augmentation technique that extends features of each graph node with aggregated information about features of its neighbors (see Appendix B.2 for a detailed discussion of NFA). Specifically, we add NFA to ResMLP and LightGBM, since they are our fastest graph-agnostic models. We denote these model versions with the -NFA suffix. We expect that if the graph structure is beneficial for the considered task, then NFA-augmented models will significantly outperform their counterparts without NFA. This indeed happens in our experiments, confirming the usefulness of the graph structure provided in our datasets. However, note that NFA provides only limited access to graph information (specifically, it provides access to aggregated information about features of 1-hop neighbors). Models that can use much more graph information are graph neural networks, which we discuss in the next paragraph.
@@
Most currently available GFMs do not support node property prediction tasks in graphs with arbitrary node features. Of those that do, we were able to find only two models with open weights: OpenGraph (Xia et al., 2024) and AnyGraph (Xia & Huang, 2024). Both OpenGraph and AnyGraph exploit the Transformer architecture and are pretrained with a link prediction objective on a mixture of different graph datasets. These methods differ in what data they can operate on. Specifically, OpenGraph only uses relational information and constructs node representations based on SVD factors of the adjacency matrix, while AnyGraph also uses the available node feature information and combines SVD factors for both the feature matrix and the adjacency matrix. Both these models were designed to be adapted to new node classification datasets without fine-tuning, using an in-context learning (ICL) setting instead. Specifically, they can perform link prediction in arbitrary graphs, and they cast any node classification task as a link prediction task where links to virtual nodes representing classes are predicted for unlabeled nodes. Further, we were able to reproduce the pretraining for one more GFM — GCOPE (Zhao et al., 2024) (weights for this model are not publicly available, but training code is). GCOPE also applies a projection to node features (e.g., based on SVD or attention mechanism), but exploits additional virtual nodes as graph coordinators to simultaneously process different graph datasets at the pretraining stage. The authors use GraphCL (You et al., 2020) or SimGRACE (Xia et al., 2022) as the pretraining objective. In contrast to OpenGraph and AnyGraph, GCOPE uses fine-tuning for adaptation to new node classification datasets. However, neither of the three models supports node regression. Further, we found that these GFMs cannot scale to large datasets.
@@ B.3 Experimental setup and hyperparameter selection details
Hyperparameter choice is extremely important for the performance of both GNNs and GBDT models. Thus, we conducted an extensive hyperparameter search on the validation set for all models. For the considered GBDT models, we ran $100$ iterations of Bayesian optimization using Optuna (Akiba et al., 2019). The specific hyperparameter distributions used for these models are provided in Table 4. Since GNNs are sensitive to different hyperparameters than GBDT models, we used a different hyperparameter search strategy for them. Specifically, we found that the learning rate and dropout probability (Srivastava et al., 2014) are the most important hyperparameters for our GNN implementations on our datasets. Thus, we ran grid search selecting the learning rate from $\{3\times 10^{-5},1\times 10^{-4},3\times 10^{-4},1\times 10^{-3},3\times 10^{-3}\}$ and dropout probability from $\{0,0.1,0.2\}$ (note that the highest learning rate of $3\times 10^{-3}$ often resulted in NaN issues, however, we still included it in our hyperparameter search, as in our preliminary experiments we found it to be beneficial for some of our dataset/model combinations). In our preliminary experiments we found that the performance of our GNNs is quite stable for a wide variety of reasonable architecture hyperparameter values (we found the use of skip-connections and layer normalization to be important for this stability). Hence, for our final experiments, we kept these hyperparameters fixed. We set these values as follows: the number of graph neighborhood aggregation blocks to $3$ and the hidden dimension to $512$. The only exceptions to this hidden dimension size were made for our largest datasets: to avoid GPU out-of-memory issues, we decreased the hidden dimension to $400$ for pokec-regions and to $200$ for web-traffic, web-fraud, and web-topics. For GNNs with attention-based graph neighborhood aggregation (GAT and GT), the number of attention heads was set to $4$. We used the Adam optimizer (Kingma & Ba, 2015) in all our GNN experiments. We trained each model for $1000$ steps and then selected the best step based on the performance on the validation set.
@@ Appendix D Limitations and broader impact
The aim of our benchmark is to introduce a diverse set of graph datasets for node property prediction that covers a wide range of domains and graph structural properties, including those not encountered in commonly used datasets for GML model evaluation. However, data that can be naturally represented as graphs is so widespread across different domains that no benchmark can cover them all. Thus, our collection of 14 datasets still only covers a small part of the wide range of situations where modeling data as a graph can be useful. But we hope that it will encourage the GML research community to use more diverse sets of datasets and focus on practically relevant applications where graph-structured data appears.
Our benchmark includes datasets with realistic tasks such as fraud detection and user engagement prediction. Poorly performing machine learning models used for these tasks in real-world services can negatively affect the users of these services. For example, type I errors of fraud detection systems, i.e., wrongly predicting that an innocent person is fraudulent, have an undesirable negative impact. Thus, particular care should be taken to minimize the probability of such errors. We believe that the release of high-quality and properly anonymized datasets for these tasks such as the ones in our benchmark will encourage the community to develop better models, since the community will be able to use these datasets as a realistic and reliable testbed to investigate which methods lead to reductions in undesirable model errors.

### v4

@@ B.1 Models
The models discussed above are graph-agnostic. To see if they can be improved by being granted access to some of the graph information, we augment them with neighborhood feature aggregation (NFA) — a simple feature augmentation technique that extends features of each graph node with aggregated information about features of its neighbors (see Appendix B.2 for a detailed discussion of NFA). Specifically, we add NFA to ResMLP and LightGBM, since they are our fastest graph-agnostic models. We denote these model versions with the -NFA suffix. We expect that if the graph structure is beneficial for the considered task, then NFA-augmented models will significantly outperform their counterparts without NFA. This indeed happens in our experiments, confirming the usefulness of the graph structure provided in our datasets. However, note that NFA provides only limited access to graph information, specifically the aggregated features of 1-hop neighbors. Models that can use much more graph information are graph neural networks, which we discuss in the next paragraph.
@@
Most currently available GFMs do not support node property prediction tasks in graphs with arbitrary node features. Among those that do, we were able to find only two models with open pretrained checkpoints: OpenGraph (Xia et al., 2024) and AnyGraph (Xia & Huang, 2024). Both OpenGraph and AnyGraph use the Transformer architecture and are pretrained with a link prediction objective on a mixture of different graph datasets. These methods differ in what data they can operate on. Specifically, OpenGraph only uses relational information and constructs node representations based on SVD factors of the adjacency matrix, while AnyGraph also uses the available node feature information and combines SVD factors for both the feature matrix and the adjacency matrix. Both these models were designed to be adapted to new node classification datasets without fine-tuning, using an in-context learning (ICL) setting. Specifically, they can perform link prediction in arbitrary graphs, and they cast any node classification task as a link prediction task where links to virtual nodes representing classes are predicted for unlabeled nodes.
Further, we reproduced the pretraining for two more GFMs — TS-GNN (Finkelshtein et al., 2025) and GCOPE (Zhao et al., 2024) (the weights for these models are not publicly available, but the training code is). GCOPE also applies a projection to node features (e.g., based on SVD or attention mechanism), but exploits additional virtual nodes as graph coordinators to simultaneously process different graph datasets at the pretraining stage. The authors use GraphCL (You et al., 2020) or SimGRACE (Xia et al., 2022) as the pretraining objective. In contrast to the ICL approaches, GCOPE uses fine-tuning for adaptation to new node classification datasets.
The TS-GNN model takes a very different approach. Its authors observe that a GFM is expected to be feature-permutation-invariant, label-permutation-equivariant, and node-permutation-equivariant, design a linear transformation that satisfies these properties, and use it as the main GFM building block. Note that TS-GNN requires solving several least squares problems to fit its weights before making predictions, which makes it not a true in-context learning model, but, following the authors of this model, we still mark it as ICL due to these least squares problems being computationally light compared to typical fine-tuning. The original paper proposes two versions of TS-GNN: TS-Mean and TS-GAT; we use TS-Mean in all our experiments.
Note that neither of the considered GFMs supports node regression (for OpenGraph and AnyGraph this is an inherent limitation of the model design, while for TS-GNN and GCOPE this is only a limitation of the official model implementations, since in theory these two models can perform node regression). Our experiments also show that these models cannot scale to large datasets. These two issues prevent us from successfully evaluating the considered GFMs on a significant number of our datasets.
@@ B.3 Experimental setup and hyperparameter selection details
Hyperparameter choice is extremely important for the performance of both GNNs and GBDT models. Thus, we conducted an extensive hyperparameter search on the validation set for all models. For the considered GBDT models, we ran $100$ iterations of Bayesian optimization using Optuna (Akiba et al., 2019). The specific hyperparameter distributions used for these models are provided in Table 4. Since GNNs are sensitive to different hyperparameters than GBDT models, we used a different hyperparameter search strategy for them. Specifically, we found that the learning rate and dropout probability (Srivastava et al., 2014) are the most important hyperparameters for our GNN implementations on our datasets. Thus, we ran grid search selecting the learning rate from $\{3\times 10^{-5},1\times 10^{-4},3\times 10^{-4},1\times 10^{-3},3\times 10^{-3}\}$ and dropout probability from $\{0,0.1,0.2\}$ (note that the highest learning rate of $3\times 10^{-3}$ often resulted in NaN issues; however, we still included it in our hyperparameter search, as in our preliminary experiments we found it to be beneficial for some of our dataset/model combinations). In our preliminary experiments we found that the performance of our GNNs is quite stable for a wide variety of reasonable architecture hyperparameter values (we found the use of skip-connections and layer normalization to be important for this stability). Hence, for our final experiments, we kept these hyperparameters fixed. We set these values as follows: the number of graph neighborhood aggregation blocks to $3$ and the hidden dimension to $512$. The only exceptions to this hidden dimension size were made for our largest datasets: to avoid GPU out-of-memory issues, we decreased the hidden dimension to $400$ for pokec-regions and to $200$ for web-traffic, web-fraud, and web-topics. For GNNs with attention-based graph neighborhood aggregation (GAT and GT), the number of attention heads was set to $4$. We used the Adam optimizer (Kingma & Ba, 2015) in all our GNN experiments. We trained each model for $1000$ steps and then selected the best step based on the performance on the validation set.
@@ Appendix D Limitations and broader impact
The aim of our benchmark is to introduce a diverse set of graph datasets for node property prediction that covers a wide range of domains and graph structural properties, including those not encountered in commonly used datasets for GML model evaluation. However, data that can be naturally represented as graphs is so widespread across different domains that no benchmark can cover them all. Thus, our collection of 14 datasets still only covers a small part of the wide range of situations where modeling data as a graph can be useful. Nevertheless, we hope that it will encourage the GML research community to use more diverse sets of datasets and focus on practically relevant applications where graph-structured data appears.
Our benchmark includes datasets with realistic tasks such as fraud detection and user engagement prediction. Poorly performing machine learning models used for these tasks in real-world services can negatively affect the users of these services. For example, type I errors of fraud detection systems, i.e., wrongly predicting that an innocent person is fraudulent, have an undesirable negative impact. Thus, particular care should be taken to minimize the probability of such errors. We believe that the release of high-quality and properly anonymized datasets for these tasks such as those in our benchmark will encourage the community to develop better models, since the community will be able to use these datasets as a realistic and reliable testbed to investigate which methods lead to reductions in undesirable model errors.